# 1]FAIR at Meta

### v1

@@ 5 Conclusion, limitations, and future work
Prior work (Momennejad et al., 2023; Valmeekam et al., 2023a) has found that LLMs, which are based on Transformer models, struggle with solving complex decision making tasks. Searchformer demonstrates that with appropriate training data, Transformers can in fact solve complex planning tasks. Moreover, Searchformer robustly follows the intermediate steps—the execution trace—of a symbolic planner and improves (in terms of trace length) beyond the human-crafted rule-based planning strategy it was initially trained on. Compared to solution-only models that directly predict a solution, our search-augmented models require fewer training sequences and scale better to more difficulty tasks.
Limitations
@@
Currently, Searchformer is trained on the execution traces of $A^{*}$ to learn a complex planning strategy. However, the trace length may grow exponentially in the length of an optimal plan (see Figure 3), and training on the resulting token sequence data become computationally very expensive. In fact, the presented experiments use token sequences that are significantly longer than the sequences used to train LLMs such as Llama 2 (Touvron et al., 2023). [^3] The 175 million parameter model trained on the Sokoban dataset uses 64 GPUs with 32GB of memory.
Future work
@@
One way to mitigate this limitation is to use curriculum learning: starting from simple tasks with reasonably long execution traces, train and fine-tune the Searchformer to reduce the trace length, and then adapt the improved model to more complex tasks. Another possibility is to integrate better heuristics or value functions into $A^{*}$ search, similar to Monte Carlo Tree Search (MCTS) (Coulom, 2006), to cap the maximal depth the search algorithm explores. Integrating hierarchical planning methods and temporal abstractions (Sutton et al., 2023, 1999; Dietterich, 2000; Hafner et al., 2022) are another avenue. This would equip the resulting model with the ability to abstract over multiple time steps and states to find an optimal plan using fewer computation steps.
We hope that our study informs further research about how to build Transformer-based models that can robustly solve decision making tasks and mitigate the limitations previously outlined by Momennejad et al. and Valmeekam et al..
@@ 6 Broader Impact
Our work focuses on symbolic planning tasks and uses synthetic datasets for training. While the tasks we explored in this paper can be easily solved with simple symbolic solvers, it is important to study the effectiveness of neural networks on such tasks. Neural network based Large Language Models have shown impressive performance on non-symbolic tasks, such as following complex natural language instructions, but still fail on relatively simple symbolic tasks like adding numbers. While the results in this paper are encouraging, they are still far from being competitive with symbolic solvers.
@@ Acknowledgment
We would like to thank Qinqing Zheng for helpful discussions and comments on earlier versions of this paper. We would also like to thank Amy Zhang for helpful discussions on this work.
Appendix A Appendix
A.1 Training encoder-decoder Transformers to predict optimal plans
@@
A.1.1 Encoder-decoder architecture
@@
The encoder network $f_{\boldsymbol{\theta}_{\text{enc}}}$ and decoder network $g_{\boldsymbol{\theta}_{\text{dec}}}$ internally both use a number of stacked causal attention layers to form an encoder-decoder Transformer (Vaswani et al., 2017) as outlined in more detail in Appendix A.2. We denote a concatenation of all encoder parameters $\boldsymbol{\theta}_{\text{enc}}$ and decoder parameters $\boldsymbol{\theta}_{\text{dec}}$ with the vector $\boldsymbol{\theta}$.
A.1.2 Training with teacher forcing
@@ A.2 Network architecture and hyper-parameters
Parameter 15M param. model 46M param. model 175M param. model 747M param. model
@@
Table 3: Optimization hyper-parameters. Every model was optimized using AdamW (Loshchilov and Hutter, 2019) with setting $\beta_{0}=0.9$ and $\beta_{1}=0.99$. Initially, the learning rate was linearly interpolated: Starting at zero and then increasing linearly to the value listed below until step 2000. Then a cosine learning rate schedule was followed (Loshchilov and Hutter, 2016). to the between zero at the first training step and the listed value below for the
@@
A.3 Dataset generation
@@
A.4 Searchformer performance analysis
@@
Figure 12: Sequence length comparison between sequences generated with $A^{*}$ search and sequences generated with each model. Each dot in each scatter plot corresponds to one specific test task. On the $x$-axis, we plot the average token sequence length when $A^{*}$ search is used. On the $y$-axis, we plot the average token sequence length when each model is used. Error bars indicate standard deviations. Percentages indicate the fraction of the generated sequences that are included in each scatter plot. 12(a): Sequence length comparison for all test prompts for which an optimal plan was generated using the corresponding model. 12(b): Sequence length comparison for all test prompts for which a correct plan was generated using the corresponding model. This plot aggregates across sequences ending in an optimal plan and sequences ending in a correct but sub-optimal plan. 12(c): Sequence length comparison for all test prompts for which a correct but sub-optimal plan was generated using the corresponding model. This plot only aggregates across sequences ending in a correct but sub-optimal plan. Sequences ending in an optimal plan were withheld.
In Sec. 3.3, each search-augmented and Searchformer model is evaluated by generating 64 token sequences for each Sokoban test task. For the same test task the same model can generate sequences that end in an optimal plan, an incorrect plan, or a correct but sub-optimal plan. In Figure 12 we compare the length of the generated sequences with the length of sequences generated when using $A^{*}$ search for each case. The percentages in each panel’s caption list how many out of all generated sequences end in an optimal plan, a correct plan (that can be optimal or sub-optimal), and a correct but sub-optimal plan.
A.5 Supplemental maze experiments
@@
Figure 14: Optimal plan prediction performance for search-augmented models. Training an encoder-decoder Transformer to imitate A* execution traces significantly boosts performance. 14(a): Percentage of correctly generated responses. Note that the plan only models only need to generate a few hundred token long plan to answer a prompt correctly. The trace plan model must generate a much longer A* execution trace correctly to correctly answer a prompt. This requires the trace plan models to generate response sequences that are hundreds or thousands of tokens long (cf. Figure 3). If one token is incorrectly predicted, the response is counted as incorrect. 14(b): Percentage of correctly generated plans. Each model was used to generate multiple responses for each prompt. If one of the generated responses contains an optimal plan, the test prompt is counted as correctly answered.
Figure 13 and Figure 14 plot how often a correct answer was provided by each model for all studied maze shape, training dataset size, and model parameterization combinations.

### v2

@@ 5 Discussion
Prior work (Momennejad et al., 2023; Valmeekam et al., 2023a) has found that LLMs struggle with solving complex decision making tasks. Searchformer demonstrates that with appropriate training data, Transformers can in fact solve complex planning tasks. Moreover, Searchformer robustly follows the intermediate steps—the execution trace—of a symbolic planner and improves (in terms of trace length) beyond the human-crafted rule-based planning strategy it was initially trained on. Compared to solution-only models that directly predict a solution, our search-augmented models require fewer training sequences and scale better to more complex planning tasks.
5.1 Limitations
@@
Currently, Searchformer is trained on the execution traces of $A^{*}$ to learn a complex planning strategy. However, the trace length may grow exponentially in the length of an optimal plan (see Figure 7), and training on the resulting token sequence data can become computationally very expensive. In fact, the presented experiments use token sequences that are significantly longer than the sequences used to train LLMs such as Llama 2 (Touvron et al., 2023).
5.2 Future work
@@
One way to mitigate this limitation and improve the efficiency of the presented methods is to use curriculum learning: starting from simple tasks with reasonably long execution traces, train and fine-tune the Searchformer to reduce the trace length, and then adapt the improved model to more complex tasks. Another possibility is to explore other planning algorithms or integrate better heuristics or value functions into $A^{*}$ search, similar to MCTS, to cap the maximal depth the search algorithm explores. Integrating hierarchical planning methods and temporal abstractions (Sutton et al., 2023, 1999; Dietterich, 2000; Hafner et al., 2022) are another avenue. This would equip the resulting model with the ability to abstract over multiple time steps and states to find an optimal plan using fewer computation steps.
In comparison to Plansformer (Pallagani et al., 2022), the presented work demonstrates how to train Transformers from scratch to solve complex planning tasks on synthetic datasets. We believe that our results and methods could be combined with methods such as Plansformer to fine-tune LLMs and enable them to solve complex planning tasks more robustly. Ultimately, we hope that our study sheds light on how Transformers can be used for multi-step planning and we hope to inform further research about better understanding the reasoning capabilities of LLMs.
@@ 6 Broader Impact
Our work focuses on symbolic planning tasks and uses synthetic datasets for training. While the tasks we explored in this paper can be easily solved with simple symbolic solvers, it is important to study the effectiveness of neural networks on such tasks. Here, we provide a proof of concept on how Transformer-based neural networks can be used to robustly solve complex planning tasks. With our work, we hope to inform further research into better understanding the reasoning capabilities of Large Language Models.
@@ Acknowledgment
We would also like to thank Amy Zhang for helpful discussions on this work.
Appendix A Training encoder-decoder Transformers to predict optimal plans
@@
A.1 Encoder-decoder architecture
@@
The encoder network $f_{\boldsymbol{\theta}_{\text{enc}}}$ and decoder network $g_{\boldsymbol{\theta}_{\text{dec}}}$ internally both use a number of stacked causal attention layers to form an encoder-decoder Transformer (Vaswani et al., 2017) as outlined in more detail in Appendix B. We denote a concatenation of all encoder parameters $\boldsymbol{\theta}_{\text{enc}}$ and decoder parameters $\boldsymbol{\theta}_{\text{dec}}$ with the vector $\boldsymbol{\theta}$.
A.2 Training with teacher forcing
@@ Appendix B Network architecture and hyper-parameters
Parameter 15M model 46M model 175M model 747M model
@@
Table 3: Optimization hyper-parameters. Every model was optimized using AdamW (Loshchilov & Hutter, 2019) with setting $\beta_{0}=0.9$ and $\beta_{1}=0.99$. Initially, the learning rate was linearly interpolated: Starting at zero and then increasing linearly to the value listed below until step 2000. Then a cosine learning rate schedule was followed (Loshchilov & Hutter, 2016). to the between zero at the first training step and the listed value below for the
@@
Appendix C Dataset generation
Figure 5: Example Sokoban puzzle [^2] This level image was composed using image icons from https://github.com/morenod/sokoban (accessed 2023-11-21).. The prompt, $A^{*}$ execution trace, and optimal plan for this task is illustrated in  Figure 6 in Appendix C. For Sokoban, we only use a non-deterministic $A^{*}$ implementation.
@@ Appendix C Dataset generation
Appendix D Evaluation criteria
In the presented experiments we evaluate if each model outputs an optimal or feasible plan and how long the generated search sequences are for the search-augmented and Searchformer models.
D.1 Measuring plan optimality
  • Any-optimal-64 criterion. For each task, we sample 64 responses from a trained model. Each response is parsed and evaluated if it contains a feasible or optimal plan, regardless of the generated <trace> part. If any of the 64 plans is feasible and optimal, then the task is labelled as correct.
    $\text{SWC}:=\frac{1}{n}\sum_{i=1}^{n}c_{i}\frac{l_{i}^{*}}{\max\{l_{i},l_{i}^{*}\}}.$
D.2 Measuring search dynamics length
For sequences ending in an optimal or feasible plan, we evaluate the length of sequence dynamics in terms of number of tokens using one of the two criteria:
    $\text{ILR}:=\frac{1}{n}\sum_{i=1}^{n}c_{i}\frac{t_{i}^{*}}{t_{i}}.$
Appendix E Searchformer performance analysis
In Section 3.3, each search-augmented and Searchformer model is evaluated by generating 64 token sequences for each Sokoban test task. For the same test task the same model can generate sequences that end in an optimal plan, an incorrect plan, or a correct but sub-optimal plan. In Figure 10 we compare the length of the generated sequences with the length of sequences generated when using $A^{*}$ search for each case. The percentages in each panel’s caption list how many out of all generated sequences end in an optimal plan, a correct plan (that can be optimal or sub-optimal), and a correct but sub-optimal plan.
Appendix F Supplemental figures for experiments
@@
Figure 10: Sequence length comparison between sequences generated with $A^{*}$ search and sequences generated with each model. Each dot in each scatter plot corresponds to one specific test task. On the $x$-axis, we plot the average token sequence length when $A^{*}$ search is used. On the $y$-axis, we plot the average token sequence length when each model is used. Error bars indicate standard deviations. Percentages indicate the fraction of the generated sequences that are included in each scatter plot. 10(a): Sequence length comparison for all test prompts for which an optimal plan was generated. 10(b): Sequence length comparison for all test prompts for which a correct plan was generated. This plot aggregates across sequences ending in an optimal plan and sequences ending in a correct but sub-optimal plan. 10(c): Sequence length comparison for all test prompts for which a correct but sub-optimal plan was generated using the corresponding model. This plot only aggregates across sequences ending in a correct but sub-optimal plan.
@@
Figure 12: Optimal plan prediction performance for search-augmented models. Training an encoder-decoder Transformer to imitate A* execution traces significantly boosts performance. 12(a): Percentage of correctly generated responses. Note that the plan only models only need to generate a few hundred token long plan to answer a prompt correctly. The trace plan model must generate a much longer A* execution trace correctly to correctly answer a prompt. This requires the trace plan models to generate response sequences that are hundreds or thousands of tokens long (cf. Figure 7). If one token is incorrectly predicted, the response is counted as incorrect. 12(b): Percentage of correctly generated plans. Each model was used to generate multiple responses for each prompt. If one of the generated responses contains an optimal plan, the test prompt is counted as correctly answered.
Figure 13: Comparison of search dynamics length (in terms of number of tokens) between $A^{*}$ search and our models (search-augmented in blue, Searchformer step 1-3 in orange), on the test subset in which our models yield optimal plans. Here, for each test task, we average the lengths of sequences that ends in an optimal plan, out of 64 trials (i.e., average-on-optimal length (Appendix D)). The box plots show the distribution of average-on-optimal lengths, in which the left boundary, mid solid line, right boundary represents the 25%, 50% and 75% quantiles. Dotted lines are means and whiskers show the range.